# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 8.3. Toric degenerations and intrinsic mirrors
By Lemma 8.4, for $k$ large enough, there exists a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $k\mathscr{P}$ such that the corresponding quasi-projective Kulikov degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ admits a birational contraction to a toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ in the sense of [47]. A mirror construction for so-called positive and simple polarized toric degenerations of Calabi–Yau varieties has been introduced by Gross–Siebert [47] – see [39] for an exposition in dimension two. The toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is an example of positive and simple toric degeneration, since its dual intersection complex is an integral affine manifold with only focus-focus singularities localized in the interior of the edges of the polyhedral decomposition. Unlike the intrinsic mirror construction for log smooth degenerations [48, 49] reviewed in §5 and used in §6-7, which is based on the enumerative geometry of punctured Gromov–Witten invariants, the mirror construction for toric degenerations is based on a scattering diagram recursively constructed from explicit initial data, and so is mostly combinatorial. In particular, it is often easier to determine the output of the mirror construction for toric degenerations. In this section, we compare the intrinsic mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\widetilde{\mathscr{P}}})$ to $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the mirror to the toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$. Similar comparison results have been obtained in [41, §3.3], [16], [71, §6.1], [17, §2].
Let $G$ be the face of $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})$ spanned by the finitely many internal exceptional curves blow down by $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ as in Lemmas 8.3 and 8.4. Consider the monoid ideal $J_{G}=NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\setminus G$, and similarly as in [41, Definition 3.14] define the Gross–Siebert locus as the open torus $T^{gs}=\mathrm{Spec}\,\mathbb{C}[G^{\mathrm{gp}}]$ in the toric stratum $\mathrm{Spec}\,\mathbb{C}[G]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})/J_{G}]$ of the toric variety $S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$.
@@
The mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})$ is constant with fiber $Y_{0_{\mathcal{X}_{\widetilde{\mathscr{P}}}}}$ in restriction to $\mathrm{Spec}\,\mathbb{C}[G]$, and so to $T^{gs}$.
@@
The map $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ induces a natural inclusion $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\subset NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\oplus G^{\mathrm{gp}}$, and so a morphism
$\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})]\times T^{gs}\longrightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\,.$
@@
$(\mathbb{A}^{1})_{H}\subset\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})]\,.$
@@
On the other hand, the choice of $H$ naturally defines a strictly convex multivalued PL function $\varphi^{gs}_{H}$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by generically pushing away the singularities from the vertices to irrational points on edges of $\widetilde{\mathscr{P}}$ as described before Lemma 8.3. Indeed, by standard toric geometry, the ample divisor $H$ restricted to each of the toric irreducible components of $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ defines strictly convex PL functions around the various vertices of $\widetilde{\mathscr{P}}$. Moreover, the choice of the point $a\in T^{gs}$, viewed as an algebra homomorphism $a:\mathbb{C}[G^{gp}]\rightarrow\mathbb{C}$, naturally defines functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ attached to the monodromy invariant directions ±miplus-or-$\pm m_{i}$ of the focus-focus singularity associated to the contracted curve $E_{i}$. Taking as input the integral affine $(kP,\widetilde{\mathscr{P}})^{gs}$, having only focus-focus singularities placed at irrational points on the edges of $\widetilde{\mathscr{P}}$, the strictly convex multivalued PL function $\varphi^{gs}_{H}$, and the functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ as “slab functions”, the mirror construction for toric degeneration of [47] – see also [39, Chapter 6], produces a projective family $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$, with a divisor $\hat{\mathcal{D}}_{H,a}^{gs}$, and an ample line bundle ^$\hat{\mathcal{L}}^{gs}_{H,a}$.
@@
The mirror family of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in a scattering diagram $\mathfrak{D}^{gs}$ on $(kP,\widetilde{\mathscr{P}})^{gs}$, constructed by the Kontsevich-Soibelman algorithm from explicit initial walls coming out of the focus-focus singularities – see [39, §6.3] for the construction of the scattering diagram and [45, Appendix A] for a discussion of theta functions. On the other hand, as reviewed in §5.3, the mirror family of $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in the canonical scattering diagram $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. Arguing as in [41, §3], replacing the use of [46], which applies to a single log Calabi–Yau surface, by [40, §8], which applies to a normal crossing union of log Calabi–Yau surfaces such as $\mathcal{X}_{\widetilde{\mathscr{P}},0}$, one shows that $\mathfrak{D}^{gs}$ can be naturally realized as a deformation of $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ obtained by pushing the focus-focus singularities away from the vertices of the triangulation $\widetilde{\mathscr{P}}$. Moreover, as in [41, Lemma 3.28], there is a natural one-to-one correspondence between broken lines and so a natural identification of the corresponding theta functions. The convergence of the algebra of theta functions for $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ then follows from the convergence of the algebra of theta functions for $\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. ∎
@@ 8.4. The general fiber of the mirror family
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface and let $(P,\mathscr{P})$ be an associated Symington polytope with a good polyhedral decomposition as in Construction 2.10. Let $\widetilde{\mathscr{P}}$ be a refinement of $k\mathscr{P}$ for large $k$ as in Lemma 8.4. Let $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ be the corresponding toric degeneration given by Lemma 8.4, and $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ the corresponding mirror family given by Theorem 8.6, with intersection complex the integral affine manifold with singularities $(kP)^{gs}$ obtained from $kP$ by pushing the focus-focus singularities away from the vertices as in §8.2.
The integral affine manifold with singularities $(kP)^{gs}$ is naturally an almost toric base as in [91]. Let $(Y^{sym},D^{sym},L^{sym})\rightarrow(kP)^{gs}$ be the corresponding almost toric fibration given by [91, Corollary 4.2]. This fibration is constructed locally as a toric momentum map fibration away from the focus-focus singularities, and using an explicit local model described in [91, §4.2] around the focus-focus singularities. In particular, $Y^{sym}$ is naturally a differentiable orbifold, with orbifold points in one-to-one correspondence with the vertices of $P$, and locally given by the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices.
@@
By the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, $(kP)^{gs}$ is a Symington polytope for $(Y,D,L^{\otimes k})$, and so, it follows from [91, Corollary 4.2] that $(Y,D,L^{\otimes k})$ admits an almost toric fibration over $(kP)^{gs}$ and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$. ∎
@@
The toric mirror degeneration $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ admits a gluing description in terms of local charts – see [39, §6.2.5] for details. We claim that it follows from this description that a general fiber $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$ by [91, Corollary 4.2]. Indeed, away from the finitely many ordinary double points corresponding to focus-focus singularities of $(kP)^{gs}$, the family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ is toric, and the torus fibration over the smooth part of $(kP)^{gs}$ is obtained by locally toric deformation of the torus fibrations on the toric irreducible components of the central fiber induced by the moment maps. Moreover, each of the ordinary double points admits a neighborhood locally described by the equation
@@
where the projection to $\mathbb{A}^{1}$ is given by the $t$-coordinate. There is an explicit torus fibration with a single nodal fiber on the corresponding general fiber of (8.5), which is studied in detail in [22, §5.1] – see also [24, Example 3.20][14, Example 1.8]. Analogously as in the proof of [24, Theorem 3.22], this fibration can be glued in a smooth way to the smooth torus fibration over the smooth locus of $(kP)^{gs}$. ∎
@@
Combining Lemma 8.7 and Lemma 8.8, we obtain that $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ are diffeomorphic as orbifolds. Moreover, the orbifold points of both $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ have Zariski neighborhoods isomorphic to the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices. Indeed, this follows for $(Y,D,L^{\otimes k})$ from the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, and for $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ from the explicit gluing construction of the mirror to a toric degenerations by gluing local models – see [39, §6.2.5].
Hence, one can torically resolve in an identical way the orbifold points of both $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ to obtain diffeomorphic smooth log Calabi–Yau surfaces $(\widetilde{Y}^{gs}_{t},\widetilde{D}^{gs}_{t},\widetilde{L}^{gs}_{t})$ and $(\widetilde{Y},\widetilde{D},\widetilde{L}^{\otimes k})$. Therefore, by [34, Theorem 5.14], the smooth log Calabi–Yau surfaces $(\widetilde{Y}^{gs}_{t},\widetilde{D}^{gs}_{t})$ and $(\widetilde{Y},\widetilde{D})$ are deformation equivalent. Moreover, for a smooth log Calabi–Yau surface, one has $H^{1}(\mathcal{O})=0$, and so algebraic line bundles are classified up to algebraic isomorphism by their first Chern classes. It follows that the line bundles $\widetilde{L}^{gs}_{t}$ and $\widetilde{L}^{\otimes k}$ are isomorphic under the deformation equivalence between $(\widetilde{Y}^{gs}_{t},\widetilde{D}^{gs}_{t})$ and $(\widetilde{Y},\widetilde{D})$. Finally, Theorem 8.9 follows by contracting back the exceptional divisors obtained while resolving the orbifold points. ∎
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface and let $(P,\mathscr{P})$ be an associated Symington polytope with a good polyhedral decomposition as in Construction 2.10. Then, for general $t$ in the open torus $T_{\mathcal{X}_{\mathscr{P}}}\subset S_{\mathcal{X}_{\mathscr{P}}}$, the fiber $(Y_{t},D_{t},L_{t})$ over $t$ of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\mathscr{P}}}$ is deformation equivalent to $(Y,D,L)$.
@@
By Theorem 8.6, there exists a fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ which is a general fiber of one of the mirror families $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$. Hence, it follows from Theorem 8.9 that the general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ is deformation equivalent to $(Y,D,L^{\otimes k})$. On the other hand, denoting by $(Y_{t},D_{t},L_{t})$ a general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, it follows from Theorem 8.2 that $(Y_{t},D_{t},L_{t}^{\otimes k})$ is a fiber of the double mirror of $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, and so is also deformation equivalent to $(Y,D,L^{\otimes k})$. Finally, as $\mathrm{Pic}(Y)$ is torsion-free by Lemma 2.9, $L_{t}^{\otimes k}\simeq L^{\otimes k}$ implies $L_{t}\simeq L$, and this concludes the proof of Theorem 8.10. ∎

### v3

@@ 8.3. Toric degenerations and intrinsic mirrors
[…]
Let $G$ be the face of $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})$ spanned by the finitely many interior exceptional curves blow down by $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ as in Lemmas 8.4 and 8.5. Consider the monoid ideal $J_{G}=NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})\setminus G$, and similarly as in [44, Definition 3.14] define the Gross–Siebert locus as the open torus $T^{gs}=\mathrm{Spec}\,\mathbb{C}[G^{\mathrm{gp}}]$ in the toric stratum $\mathrm{Spec}\,\mathbb{C}[G]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})/J_{G}]$ of the toric variety $S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$.
@@
[…]
@@
The map $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ induces a natural inclusion $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})\subset NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})\oplus G^{\mathrm{gp}}$, and so a morphism
$\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})]\times T^{gs}\longrightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\,$.
@@
$(\mathbb{A}^{1})_{H}\subset\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})]\,.$
@@
[…]
@@
The mirror family of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in a scattering diagram $\mathfrak{D}^{gs}$ on $(kP,\widetilde{\mathscr{P}})^{gs}$, constructed via the Kontsevich-Soibelman algorithm from explicit initial walls coming out of the focus-focus singularities; see [42, §6.3] for the construction of the scattering diagram and [48, Appendix A] for a discussion of theta functions. On the other hand, as reviewed in §5.3, the mirror family of $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in the canonical scattering diagram $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. Arguing as in [44, §3], replacing the use of [49], which applies to a single log Calabi–Yau surface, by [43, §8], which applies to a normal crossing union of log Calabi–Yau surfaces such as $\mathcal{X}_{\widetilde{\mathscr{P}},0}$, we obtain that $\mathfrak{D}^{gs}$ naturally arises as a deformation of $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, obtained by pushing the focus-focus singularities away from the vertices of the triangulation $\widetilde{\mathscr{P}}$. Furthermore, as in [44, Lemma 3.28], there exists a natural one-to-one correspondence between broken lines, which induces a canonical identification of the corresponding theta functions. The convergence of the algebra of theta functions for $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ then follows from the convergence of the algebra of theta functions for $\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. ∎
@@ 8.4. The general fiber of the mirror family
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and let $(P,\mathscr{P})$ be an associated Symington polytope with a good polyhedral decomposition as in Construction 2.10. Let $\widetilde{\mathscr{P}}$ be a refinement of $k\mathscr{P}$ for large $k$ as in Lemma 8.5. Let $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ be the corresponding toric degeneration given by Lemma 8.5, and $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ the corresponding mirror family given by Theorem 8.7, with intersection complex the integral affine manifold with singularities $(kP)^{gs}$ obtained from $kP$ by pushing the focus-focus singularities away from the vertices as in §8.2.
The integral affine manifold with singularities $(kP)^{gs}$ is naturally an almost toric base as in [95]. Let $(Y^{sym},D^{sym},L^{sym})\rightarrow(kP)^{gs}$ be the corresponding almost toric fibration given by [95, Corollary 4.2]. This fibration is constructed locally as a toric momentum map fibration away from the focus-focus singularities, and using an explicit local model described in [95, §4.2] around the focus-focus singularities. In particular, $Y^{sym}$ is naturally a differentiable orbifold, with orbifold points in one-to-one correspondence with the vertices of $P$, and locally given by the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices.
@@
By the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, $(kP)^{gs}$ is a Symington polytope for $(Y,D,L^{\otimes k})$. Thus, by [95, Corollary 4.2], $(Y,D,L^{\otimes k})$ admits an almost toric fibration over $(kP)^{gs}$ and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$. ∎
@@
The toric mirror degeneration $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ admits a gluing description in terms of local charts – see [42, §6.2.5] for details. We claim that it follows from this description that a general fiber $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$ by [95, Corollary 4.2]. Indeed, away from the finitely many ordinary double points corresponding to focus-focus singularities of $(kP)^{gs}$, the family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ is toric, and the torus fibration over the smooth part of $(kP)^{gs}$ is obtained by locally toric deformation of the torus fibrations on the toric irreducible components of the central fiber induced by the moment maps. Moreover, each of the ordinary double points admits a neighborhood locally described by the equation
@@
where the projection to $\mathbb{A}^{1}$ is given by the $t$-coordinate. There is an explicit torus fibration with a single nodal fiber on the corresponding general fiber of (8.5), which is studied in detail in [22, §5.1] – see also [25, Example 3.20][14, Example 1.8]. Analogously as in the proof of [25, Theorem 3.22], this fibration can be glued in a smooth way to the smooth torus fibration over the smooth locus of $(kP)^{gs}$. ∎
@@
Combining Lemma 8.8 and Lemma 8.9, we obtain that $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ are diffeomorphic as orbifolds. Furthermore, the orbifold points of both $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ have Zariski neighborhoods isomorphic to toric surface singularities with momentum polytopes given by the cones describing $P$ near its vertices. Indeed, this follows for $(Y,D,L^{\otimes k})$ from the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, and for $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ from the explicit gluing construction of the mirror to a toric degenerations by gluing local models – see [42, §6.2.5].
[…]
@@
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and let $(P,\mathscr{P})$ be an associated Symington polytope with a good polyhedral decomposition as in Construction 2.10. Then, for general $t$ in the open torus $T_{\mathcal{X}_{\mathscr{P}}}\subset S_{\mathcal{X}_{\mathscr{P}}}$, the fiber $(Y_{t},D_{t},L_{t})$ over $t$ of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\mathscr{P}}}$ is deformation equivalent to $(Y,D,L)$.
@@
By Theorem 8.7, there exists a fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ which is a general fiber of one of the mirror families $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$. Hence, by Theorem 8.10, the general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ is deformation equivalent to $(Y,D,L^{\otimes k})$. On the other hand, denoting by $(Y_{t},D_{t},L_{t})$ a general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, it follows from Theorem 8.2 that $(Y_{t},D_{t},L_{t}^{\otimes k})$ is a fiber of the double mirror of $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, and so is also deformation equivalent to $(Y,D,L^{\otimes k})$. Finally, as $\mathrm{Pic}(Y)$ is torsion-free by Lemma 2.9, $L_{t}^{\otimes k}\simeq L^{\otimes k}$ implies $L_{t}\simeq L$, and this concludes the proof of Theorem 8.11. ∎